OpenStudy (pink.1d):

What sine function represents an amplitude of 4, a period of pi/2 , no horizontal shift, and a vertical shift of -3?

OpenStudy (pink.1d):

PLEASE HELP ASAP I WILL FAN AND MEDAL :)

OpenStudy (pink.1d):

@triciaal @Hero @Preetha

OpenStudy (pink.1d):

@Kainui @dan815

OpenStudy (pink.1d):

The choices are f(x) = -3 sin 4x + 4 f(x) = 4 sin 4x - 3 f(x) = 4 sin x - 3 f(x) = -3 sin x + 4

Vocaloid (vocaloid):

@pink.1d The general form of a sin function is a*sin(bx-c) + d a = the amplitude, which is 4 in this case the period is 2pi/b d is the vertical shift (-3 in this case)

OpenStudy (pink.1d):

@Vocaloid so the answer is 4 sin something x -3?

OpenStudy (pink.1d):

@Sepeario

OpenStudy (pink.1d):

@Vocaloid what is the bx-c PLEASEEEEEEEE

OpenStudy (pink.1d):

@Vocaloid @Vocaloid

OpenStudy (pink.1d):

@Sepeario ok i c u viewing my question for a long time so is there any way u can help?

OpenStudy (pink.1d):

@freckles

OpenStudy (sepeario):

I'm sorry I don't really work with this kind of stuff.. :(

OpenStudy (pink.1d):

Oh its ok im really desperate at this poniinnntt

OpenStudy (pink.1d):

@Vocaloid omg finally u came back to my quesstion

OpenStudy (pink.1d):

so the answer is 4 sin something x -3? what is bx-c

Vocaloid (vocaloid):

@pink.1d c is the horizontal shift, which is 0 in this case you can find b using the fact that the period = 2pi/b

OpenStudy (pink.1d):

@Vocaloid what is the answer tell me

OpenStudy (pink.1d):

i will medal and fan PLEASE

Vocaloid (vocaloid):

@pink.1d: I've given you everything you need to know period = pi/2 = 2pi/b, you can now solve for b

OpenStudy (pink.1d):

what is b?

OpenStudy (pink.1d):

ya u r

OpenStudy (pink.1d):

i do that

OpenStudy (pink.1d):

hahaha

Vocaloid (vocaloid):

@pink.1d: Please, all you need to do is solve pi/2 = 2pi/b

OpenStudy (pink.1d):

oh ok ok ok fine

OpenStudy (pink.1d):

ITS 4

Vocaloid (vocaloid):

@pink.1d: see, that wasn't so hard, was it? :)

OpenStudy (pink.1d):

f(x) = 4 sin 4x - 3

OpenStudy (pink.1d):

THATS THE ANSWER RIGHT

OpenStudy (pink.1d):

i thought it was something else thats y i didnt want to solve cuz i thought i didnt know how to but i got mixed w something else

Vocaloid (vocaloid):

@pink.1d Yup! I wanted you to find b yourself so you would understand how to do it, instead of just me explicitly giving you the answer. Everyone wins that way ~

OpenStudy (pink.1d):

haha ur smart and THANK YOU for once i understand something